# AsymDex: Leveraging Asymmetry and Relative Motion in Learning Bimanual Dexterity

### v1

@@ 19      Update the Policy $\pi_{\mathrm{AsymDex}}$ based on $\mathcal{B}$;
In addition to asymmetry, a key characteristic of bimanual dexterous manipulation is the synchronized motion of the two hands in which each hand moves in response to the other. In this section, we explain how we can further reduce the size of the observation and action spaces by defining relative and object-centric spaces that capture the relationships between the motions of two hands and the object(s) being manipulated. Indeed, the use of relative state spaces has shown to considerably benefit bimanual manipulation with simple end effectors [42, 43, 44, 23]. Furthermore, some of these prior work is limiting the relative space in a one-degree-of-freedom (1-DoF) action space [23], while AsymDex allows for complete 6-DoFs relative space.
Let $o_{f}=o_{1}$ be the state of the object being held by the facilitating hand, and let $o_{d}=o_{2}$ be the state of the object being manipulated by the dominant hand. We attach a coordinate frame to the object being held by the facilitating hand: $P_{f}$. Now, we can transform the observations $(\xi^{b}_{d}(t),\xi^{b}_{f}(t),\xi^{h}_{d}(t),o_{f}(t),o_{d}(t))$, which where originally defined in the world coordinate frame $P_{W}$, into the new coordinate frame $P_{f}$. Note that since there is no relative motion between the facilitating hand and the object that it’s holding, neither $o_{f}(t)$ nor $\xi^{b}_{f}(t)$ do not change in $P_{f}$. As such, both $o_{f}(t)$ and $\xi^{b}_{f}(t)$ can be neglected without losing any information. Now, transforming the remaining observations $(\xi^{b}_{d}(t),\xi^{h}_{d}(t),o_{d}(t))$ into the new coordinate $P_{f}$ yields $(\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$, respectively. Here, $\xi^{b}_{r}(t)$ denotes the 6D pose of the dominant hand base and $o_{r}(t)$ denotes the 6D pose of the object being manipulated by the dominant hand, both now defined relative to the object being held by the facilitating hand. Note that since $\xi^{h}_{d}(t)$ denotes the dominant hands’ joint states, it is not impacted by the change of coordinates. Similarly, we apply the same modifications to the action spaces of the asymmetry bimanual policy. This allows us to reduce the actions from $(\hat{\xi}^{b}_{d}(t),\hat{\xi}^{b}_{f}(t),\hat{\xi}^{h}_{d}(t))$ to $(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t))$, where $\hat{\xi}^{b}_{r}(t)$ is the target relative pose of the dominant hand now defined relative to the object being held by the facilitating hand.
Incorporating the above change of coordinates in addition to leveraging asymmetry, allows us to define AsymDex’s policy as $\pi_{\mathrm{AsymDex}}(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t)|\xi^{b}_{r}(t),\xi^{h}_{d}(t),o_{r}(t))$. Note that our formulation has significantly reduced the dimensions of both the state and action spaces, compared to the naive policy $\pi_{\mathrm{naive}}$ as defined in Section 3.2. See Appendix. A for RL algorithm and policy architecture.
3.5 Relative Pose Controller
where $R^{o_{f}}_{world}$ denotes the rotational transformation from Frame $P_{f}$ to the world frame $P_{W}$, $dist(\cdot)$ denotes the difference between two 6D poses. and $\alpha$ is a hyperparameter that controls the involvement of each hand. Throughout our experiments, we used fixed $\alpha=0.5$. The pseudo-code of the training process is included in Alg. 1.
3.6 Acquiring Objects
Figure 2: We decompose asymmetric bimanual dexterous manipulation into two phases: An acquisition phase and an interaction phase. We show that AsymDex can be integrated with learned grasping policies in order to seamlessly execute both phases.
While our approach as explained thus far deals with the challenge of coordinating two hands to accomplish asymmetric dexterous manipulation tasks, it assumes the task begins with that the object(s) of interest have already been grasped. However, in practice, robots must be able to tackle the challenge of grasping the necessary objects before the interaction between the two hands and the objects can begin. We refer to this phase as the acquisition phase. Most recent works on bimanual dexterous manipulation often entirely ignore the acquisition phase and focus purely on the interaction phase [7, 5, 40]. In contrast, we demonstrate that our approach can seamlessly accommodate the acquisition phase by i) leveraging the observation that the acquisition phase doesn’t require the coordination of two arms, and ii) employing recent advances in learning to grasp. Specifically, we demonstrate that we can seamlessly integrate AsymDex with PDGM [45], which can efficiently learn multi-fingered grasping policies by leveraging pre-grasp poses (see Fig. 2). Details about the grasping reward design are available in Appendix. B. We begin by executing the grasping policy in isolation and then ”turn on” the asymmetric policy learned by AsymDex after the object has been firmly grasped by the facilitating hand. If the task requires the dominant hand to also grasp a second object, we employ the same method to train a grasping policy for dominant hand to acquire the object, but switch the control of the dominant hand’s joints over the asymmetric policy after the object has been grasped.
4 Evaluation
We evaluated AsymDex on four asymmetric bimanual manipulation tasks and compared its performance against strong baselines that challenge our key design choices. We begin by explaining the various aspects of our experimental setup and follow with a discussion of results.
Platform: We conducted all our experiments using two ShadowHands – each a 30-DoF simulated multi-finger hand system (24-DoF hand + 6-DoF floating wrist base) built with Isaac Gym [46].
Tasks: We evaluated AsymDex and the baselines on the following four biannual manipulation tasks which contain both original (Block in cup, Bottle cap) and adapted tasks (Stack, Switch) from BiDexHand [20] (see Fig. 3 for visualization).
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
Note that the hands have to coordinate and perform all four tasks without relying on a support surface such as a table and ensure that the object(s) are not dropped. See Appendix. C for details on state space design, sampling procedure, success criteria, and reward design for each task.
Figure 3: We evaluate AsymDex on four bimanual dexterous manipulation tasks.
Metrics: We quantify performance in terms of i) Task success rate (see Appendix. C for criteria) and ii) sample efficiency. We report both metrics across five random seeds in all experiments.
We evaluated AsymDex with two sets of experiments as described below.
4.1 Learning Bimanual Coordination
In this experiment, we focus on AsymDex’s effectiveness during the interaction phase. We initialize the environment such that the hands are at a pre-grasp pose around the objects using appropriate initial grasps (After the first timestep of the environment, the hand need to learn to catch and grasp the object from pre-grasp pose if it is not a facilitating hand). Note that this is a common assumption in recent methods that learn bimanual dexterous manipulation skills [5, 40]. Further, this allows us to isolate and examine AsymDex’s ability to learn to coordinate two multi-fingered hands. See Section 4.2 for the second experiment in which we also consider the challenge of acquiring the objects from a tabletop surface before interaction begins.
We compare AsymDex policy against the following baselines:
  • $\bullet$
    Monolithic: This policy doesn’t make assumptions about the structure of bimanual manipulation (see Sec. 3.2). As such, this baseline allows us to examine the necessity and effectiveness of leveraging both the asymmetry in hand roles and the relative action and observation spaces.
  • $\bullet$
  • $\bullet$
We report the learning curves in Fig. 4 and success rates in Table 1. As can be readily observed, AsymDex is the only method that consistently performs well across all tasks, either performing comparably or outperforming all the baselines. While Rel-w/o-asym performs comparably to AsymDex on Stack, it is not able to match AsymDex’s performance on the other tasks. Both AsymDex and Rel-w/o-asym perform substantially better than the other two baselines across all tasks except Switch, where AsymDex performs much better than all three baselines. In fact, Asym-w/o-rel is able to learn an effective policy in only one task (Bottle cap), while Monolithic struggles on all four tasks. Curiously, Monolithic outperforms asym-w/o-rel on the Stack task. Qualitative analysis of rollouts reveals that, unlike Asym-w/o-rel, the Monolithic policy learns to use the facilitating hand’s fingers to orient the cup towards the bottom of the other cup.
Table 1: Success rates (mean $\pm$ std. dev.) in Interaction phase
Task Method Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
Block in cup $0.0429\pm 0.0266$ $0.0164\pm 0.0190$ $0.3052\pm 0.1577$ 0.7701 $\pm$ 0.0559
Stack $0.2517\pm 0.2809$ $0.0044\pm 0.0037$ 0.9443 $\pm$ 0.0136 $0.9233\pm 0.0250$
Bottle cap $0.6125\pm 0.1734$ $0.7893\pm 0.1277$ $0.8073\pm 0.1389$ 0.8301 $\pm$ 0.1797
Taken together, the above observations reveal a few insights. First, when used in isolation, neither asymmetry nor relative motion are sufficient across all tasks. In particular, while each might provide sufficient structure to accomplish some tasks, they prove to be less effective on other tasks. Second, the use of relative motion offers a larger boost in performance compared to asymmetry, likely due to the fact that relative spaces avoid unnecessary exploration (e.g., when the two hands move in parallel) while allowing the facilitating hand to exhibit more complex behaviors. Third, ignoring both asymmetry and relative motion hardly leads to success.
4.2 Learning to Grasp and Coordinate
In this experiment, we evaluate AsymDex’s ability to incorporate the object acquisition phase in addition to the interaction phase. Specifically, we initialize the environment for each task such that the objects of interest are placed on a tabletop surface. As such, each method needs to learn both to grasp the necessary objects and to coordinate the two hands to complete the tasks.
For AsymDex, we follow the same strategy as introduced in Section 3.6. We compare AsymDex’s performance against the following baselines:
  • $\bullet$
    1-stage-monolithic: This baseline uses a single policy to learn both the grasping and interaction phases for both hands, and thus allows us to investigate the benefits of AsymDex’s two-phase decomposition.
  • $\bullet$
    2-stage-monolithic: This policy benefits from the same phase decomposition as AsymDex policy, but leverages neither asymmetry nor relative motion. As such, this baseline allows us to examine if this task can be solved merely with two-phase decomposition.
We report the overall roll-out success rate of all methods for two tasks across five random seeds in Table. 2. We find that AsymDex significantly outperforms the other two baselines in both tasks, suggesting that combining phase decomposition with AsymDex’s other two design choices (asymmetry and relative state) results in policies that can effectively handle both the acquisition and the interaction phases of bimanual dexterous manipulation. The fact that 2-stage-monolithic baseline outperforms the 1-stage-monolithic baseline demonstrates the inherent benefits of phase decomposition. Our qualitative analysis of Block in cup task revealed that 1-stage-monolithic policy learns to tip the cup over and push the block towards the cup. In contrast, both the two-stage policies learn the expected behavior. This suggests that the phase decomposition nudges the grasping and interaction policies to learn reasonable behaviors that complement each other.
Task Method 1-stage-monolithic 2-stage-monolithic 2-stage-AsymDex
Block in cup $0.0321\pm 0.0251$ $0.1505\pm 0.1059$ 0.7938 $\pm$ 0.0897
Bottle cap $0.1680\pm 0.2695$ $0.6407\pm 0.1141$ 0.8726 $\pm$ 0.0600
5 Conclusion
6 Limitations and Future Works
Our work has revealed a number of limitations and avenues for future research. First, AsymDex in its current form cannot handle certain bimanual tasks that require complex multi-finger manipulation from both hands (e.g., reorienting a heavy object, dynamic handover). Second, AsymDex does not consider the kinodynamic constraints that might result from manipulator arms. Third, AsymDex has not yet been evaluated on hardware. Fourth, behaviors produced by AsymDex are not always natural or human-like due to lack of necessary incentives.
Appendices
Appendix A RL Training
We use Proximal Policy Optimization (PPO) [47] algorithm to train all policies $\pi_{\mathrm{naive}},\pi_{\mathrm{asym}},$ and $\pi_{\mathrm{AsymDex}}$ with their corresponding value functions. Both policies and value functions are parameterized via a three-layer MLP network. The size of hidden layers for each is i) policy: (256, 256, 128), ii) value function: (512, 512, 512). The activation functions are all set as Exponential Linear Unit (ELU). We use the same PPO hyperparameters for all the baselines and AsymDex ($learning\ rate:3\times 10^{-4},\gamma:0.98,\lambda:0.95$). We train the polices on a computer with a single Nividia RTX 4090 GPU.

### v2

@@ 4.2 Incorporating Relative Observation and Action Spaces
Figure 2: AsymDex’s observation and Action Spaces.
In addition to asymmetry, a key characteristic of bimanual dexterity is the synchronized and responsive movement of the two hands. We can further reduce the size of the observation and action spaces and introduce tight coupling between the hands’ behaviors by defining relative and object-centric coordinates that capture the relationships between the movements of two hands and the object(s) being manipulated. Indeed, the use of relative state spaces has shown to considerably benefit bimanual manipulation with simple end effectors [26, 46, 47, 48]. While some of these prior works limit the relative space to a one-degree-of-freedom (1-DoF) action space [26], AsymDex allows for complete 6-DoFs relative space.
Let $\xi^{obj}_{f}(t)$ be the state of the object being held by the facilitating hand, and let $\xi^{obj}_{d}(t)$ be the state of the object being manipulated by the dominant hand. We attach a coordinate frame to the object being held by the facilitating hand: $P_{f}$. Now, we can transform the asymmetry-only observations $z_{\mathrm{asym}}(t)$ (originally defined in the absolute or world coordinate frame) into the new coordinate frame $P_{f}$. Note that since there is no relative motion between the facilitating hand and the object that it is holding, neither $\xi^{obj}_{f}(t)$ nor $\xi^{b}_{f}(t)$ change in $P_{f}$, and can thus be removed from our observation space without losing any information. Now, transforming the observations ($\xi^{b}_{d}(t),\xi^{obj}_{d}(t)$) onto $P_{f}$ yields $z_{\mathrm{AsymDex}}=[\xi^{h}_{d}(t),\xi^{b}_{r}(t),\xi^{obj}_{r}(t)]$, where $\xi^{b}_{r}(t)$ and $\xi^{obj}_{r}(t)$ respective denote the 6D relative poses of the dominant hand base and the object being manipulated by the dominant hand, both defined with respect to the object being held by the facilitating hand. Note that since $\xi^{h}_{d}(t)$ denotes the dominant hands’ joint states, it is not impacted by the change of coordinates. Similarly, we apply the same modifications to the asymmetry-only action $a_{asym}(t)$, yielding $a_{\mathrm{AsymDex}}(t)=(\hat{\xi}^{b}_{r}(t),\hat{\xi}^{h}_{d}(t))$, where $\hat{\xi}^{b}_{r}(t)$ is the target relative pose of the dominant hand now defined relative to the object being held by the facilitating hand. Incorporating the above change of coordinates in addition to leveraging asymmetry, allows us to define AsymDex’s policy as $\pi_{\mathrm{AsymDex}}(a_{\mathrm{AsymDex}}(t)|z_{\mathrm{AsymDex}}(t))$. Note that our formulation has significantly reduced the dimensions of both the state and action spaces, compared to the symmetric policy $\pi_{\mathrm{sym}}$ as defined in Section 3.
We parameterize the AsymDex policy $\pi_{\mathrm{AsymDex}}(\cdot)$ using an MLP and using Proximal Policy Optimization (PPO) [49] to train it. See Appendix. A for details of the algorithm and policy architecture.
4.3 Relative Pose Tracking Controller
where $R^{o_{f}}_{world}$ denotes the rotational transformation from Frame $P_{f}$ to the world frame $P_{W}$, $dist(\cdot)$ denotes the difference between two 6D poses, and $\alpha$ is a hyperparameter that controls the relative involvement of each hand. The pseudo-code of the training process is included in Alg. 1.
4.4 Acquisition Phase
Figure 3: We decompose asymmetric bimanual dexterous manipulation into two phases: An acquisition phase and an interaction phase. We show that AsymDex can be readily integrated with learned grasping policies in order to seamlessly acquire and manipulate objects.
While our approach as explained thus far deals with the challenge of coordinating two hands to accomplish asymmetric dexterous manipulation tasks, it assumes that the object(s) of interest have already been grasped at the beginning of the task. However, in practice, robots must be learn to grasp and pick up the necessary objects before the interaction between the two hands (and the objects) can begin. We refer to this initial phase as the acquisition phase. Most recent works on bimanual dexterous manipulation often entirely ignore the acquisition phase and focus purely on the interaction phase [6, 19, 20]. In contrast, we demonstrate that our approach can seamlessly accommodate the acquisition phase by i) leveraging the observation that the acquisition phase doesn’t require the coordination of two arms, and ii) employing recent advances in learning to grasp. Specifically, we demonstrate that we can seamlessly integrate AsymDex with PDGM [50], which can efficiently learn multi-fingered grasping policies by leveraging pre-grasp poses (see Fig. 3). Details about the grasping reward design are available in Appendix. B. We begin by executing the grasping policy in isolation and then ”turn on” the asymmetric policy learned by AsymDex after the object has been firmly grasped by the facilitating hand. If the task requires the dominant hand to also grasp a second object, we employ the same method to train a grasping policy for dominant hand to acquire the object, but switch the control of the dominant hand’s joints over the asymmetric policy after the object has been grasped.
5 Experimental evaluation
We evaluated AsymDex on four simulated and three real-world bimanual dexterous tasks and compared its performance against strong baselines that challenge our key design choices.
5.1 Simulation Experiments
Our experiments in simulation both systematically and rigorously evaluate AsymDex.
Tasks: We evaluated AsymDex and the baselines on the following four bimanual manipulation tasks which contain both original (Block in cup, Bottle cap) and adapted tasks (Stack, Switch) from BiDexHand [14] (see Fig. 1). All these tasks use two Shadow Hands – each a 30-DoF simulated multi-finger hand system (24-DoF hand + 6-DoF floating wrist base) built with Isaac Gym [51].
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
Note that our task designs are more challenging than their counterparts in BiDexHand [14]. We require that the two hands coordinate and synchronize to achieve success in each of the above four tasks, especially since (unlike the original designs) we do not provide a support surface (e.g., a table) that would significantly reduce the need for bimanual coordination. See Appendix. C for details on state space design, sampling procedure, success criteria, and reward design.
Metrics: We quantify performance in terms of i) success rate (see Appendix. C for criteria) and ii) sample efficiency. We report both metrics across five random seeds in all experiments.
5.1.1 Learning Bimanual Coordination
We first evaluated AsymDex’s effectiveness during the interaction phase. Following contemporary practice in methods that learn bimanual dexterous skills [19, 20], we initialized the environment such that the hands are at appropriate pre-grasp poses near the respective objects. This allows us to isolate and examine AsymDex’s ability to learn to coordinate two multi-fingered hands. See Section 5.1.2 for the second experiment in which we also consider the challenge of acquiring the objects from a tabletop surface before interaction begins.
We compared AsymDex against the following baselines:
  • $\bullet$
    Sym: This policy assumes that both hands play an equal role in bimanual manipulation (see $\pi_{sym}$ in Sec. 3). This baseline allows us to examine the necessity and effectiveness of leveraging the asymmetry in hand roles as well as the relative action and observation spaces.
  • $\bullet$
  • $\bullet$
Both asymmetry and relative spaces are necessary for consistent performance: We report the learning curves in Fig. 4 (a) and success rates in Table 1. Note that AsymDex consistently outperforms all the baselines across all four tasks in terms of success rate and sample efficiency, with significant margins in two tasks (Block in cup and Switch). Rel-w/o-asym performs better than the other two baselines across all tasks except Switch, while Asym-w/o-rel performs better than Monolithic on all tasks except Block in cup, on which both struggle.
Taken together, the above observations reveal a few key insights. First, when used in isolation, neither asymmetry nor relative spaces are sufficient across all tasks. Second, the use of relative spaces offers a larger boost in performance compared to asymmetry, likely due to the fact that relative spaces avoid unnecessary exploration (e.g., when the two hands move in parallel) while allowing the facilitating hand to exhibit more complex behaviors. Third, ignoring both asymmetry and relative spaces (Sym) hardly leads to success.
Task Method Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
Block in cup $0.0429\pm 0.0266$ $0.0164\pm 0.0190$ $0.1086\pm 0.1378$ 0.7701 $\pm$ 0.0559
Stack $0.0771\pm 0.0611$ $0.2185\pm 0.3232$ $0.6560\pm 0.3213$ 0.8392 $\pm$ 0.0596
Bottle cap $0.3111\pm 0.1813$ $0.4143\pm 0.1410$ $0.4730\pm 0.2011$ 0.6295 $\pm$ 0.1422
Table 1: Success rates (mean $\pm$ std. dev.) for the interaction phase.
5.1.2 Learning to Grasp and Coordinate
We next evaluated AsymDex’s ability to incorporate the object acquisition phase before the interaction phase. Specifically, we initialize the environment for each task such that the objects of interest are placed on a tabletop surface. As such, each method needs to learn both to grasp the necessary objects and to coordinate the two hands to complete the tasks.
Task Method Monolithic 2-stage-sym AsymDex
Block in cup $0.0321\pm 0.0251$ $0.1505\pm 0.1059$ 0.7938 $\pm$ 0.0897
Bottle cap $0.0\pm 0$ 0.0 $0.2552\pm 0.1573$ 0.6116 $\pm$ 0.1328
For AsymDex, we follow the same strategy introduced in Section 4.4, and comapare its performance against the following baselines:
  • $\bullet$
    Monolithic: This baseline uses a single policy to learn both the grasping and interaction phases for both hands, allowing us to investigate the benefits of two-phase decomposition.
  • $\bullet$
    2-stage-sym: This policy benefits from the two phase decomposition but leverages neither asymmetry nor relative spaces. As such, this baseline allows us to examine if merely employing two-phase decomposition is sufficient.
AsymDex can effectively combine the acquisition and interaction phases: We report the overall roll-out success rate of all methods for two tasks across five random seeds in Table. 2. We find that AsymDex significantly outperforms the other two baselines in both tasks, suggesting that combining phase decomposition with AsymDex’s other two design choices (asymmetry and relative spaces) results in policies that can effectively handle both the acquisition and the interaction phases of bimanual dexterous manipulation. The fact that 2-stage-sym baseline outperforms the Monolithic baseline points to the inherent benefits of phase decomposition. Our qualitative analysis of Block in cup task revealed that Monolithic policy learns to tip the cup over and push the block towards the cup. In contrast, both the two-stage policies learn more intuitive behaviors, suggesting that the phase decomposition nudges the grasping and interaction policies to learn reasonable behaviors that complement each other.
5.2 Real-world Experiments
We finally evaluated AsymDex and the same baselines on the following 3 real-world bimanual manipulation tasks inspired by recent bimanual manipulation works [5, 19, 26, 52, 53, 54, 55].
Hardware setup: We use one 16-dof Allegro hand from Wonik Robotics and one 6-DoF Ability Hand from Psyonic. They are mounted on two 7-DoF Kinova Gen3 robotic arms. We used the Ability Hand for one of the hands since we only had access to one physical Allegro hand. For object pose information, we either employ an Realsense camera with AprilTag-based object tracking method or estimate it directly from the end-effector’s pose, especially during occlusion.
Real-world Tasks: i) Block in Cup: Place a block with sides of 5 cm inside a cylindrical cup with an internal diameter of 10 cm, ii) Pour: Pour dry beans from one cup into another, and iii) Twist Lid: Twist the lid off a jar while holding the jar. Snapshots are shown in Fig. 5.
Simulation Evaluation: For each task, we first construct a simulated counterpart of our hardware setup and then train all policies in simulation. We then compared their performance in terms of success rate and sample efficiency, as reported in Fig. 4 (b). As observed in previous experiments, AsymDex consistently and significantly outperformed all baselines in terms of both metrics.
Sim2Real Transfer: We deployed only the AsymDex policies on hardware since the baseline policies tended to exhibit either negligible success rates or aggressive behaviors that could damage the hardware. To enable effective sim-to-real transfer, we apply domain randomization during reinforcement learning, and details of the randomization process are provided in Appendix. D.
Block in cup Pouring Twist Lid
16/20 17/20 18/20
Table 3: Real-world task success rates for AsymDex.
AsymDex enables zero-shot Sim2Real Transfer: As shown in Table. 3, AsymDex achieves high success rates across all deployments. This once again highlights the effectiveness of AsymDex in learning robust and reliable real-world bimanual dexterous manipulation skills. For the block in cup and pour task, AsymDex is simply trained with observation noise component of common domain randomization technique to learn a policy that can be deployed in real world successfully. For the most challenging twist lid task, we added observation noise and action noise, as well as other randomization components for sim2real transfer. Though we utilized similar reward design as previous work [19] for this task, we observed that AsymDex learns a more natural and human-like behavior compared to the previous work. Besides, unlike previous work which fixes the hand base motion [2, 19], AsymDex is able to reposition and reorient both hand bases while two hands interact, enabling more fluent bimanual coordination. See Appendix. D for the detailed reward design and simulation success criterion for each task.
Figure 5: We created simulation environments to match our hardware setup.
6 Conclusion
7 Limitations and Future Work
Our work has revealed a number of limitations and avenues for future research. First, AsymDex in its current form cannot handle certain bimanual tasks that require complex multi-finger manipulation from both hands (e.g., reorienting a heavy object, dynamic handover). Second, AsymDex does not consider the kinodynamic constraints that might result from manipulator arms. Three, behaviors produced by AsymDex are not always natural or human-like due to lack of necessary incentives.
Appendices
Appendix A Training Details
We use Proximal Policy Optimization (PPO) [49] algorithm to train all policies $\pi_{\mathrm{naive}},\pi_{\mathrm{asym}},$ and $\pi_{\mathrm{AsymDex}}$ with their corresponding value functions. Both policies and value functions are parameterized via a three-layer MLP network. The size of hidden layers for each is i) policy: (256, 256, 128), ii) value function: (512, 512, 512). The activation functions are all set as Exponential Linear Unit (ELU). We use the same PPO hyperparameters for all the baselines and AsymDex ($\gamma:0.98,\lambda:0.95$, clip range: 0.2, minibatch size: 8092). We use an adaptive learning rate with KL threshold of 0.016. We train the polices on a computer with a single Nividia RTX 4090 GPU.